OpenStudy (anonymous):

I need help with this You want to decorate the side of a cylindrical can by coloring a rectangular piece of paper and wrapping it around the can. The paper is 21 cm by 32 cm. Find the two possible diameters of the can to the nearest 0.01 cm. Assume the paper fits exactly.

OpenStudy (anonymous):

anyone anyone??

OpenStudy (raden):

one of the measure of paper can be a diameter for size 21, it mean the perimeter of base of cylindrical = 21 so, pi * d = 21 solve for d d = 21/pi = 21/(3.14) = ... do same idea for size 32

OpenStudy (anonymous):

Wow thxs that makes alot of sense now... THXS:)

OpenStudy (raden):

you're welcome :)